One canned juice drink is 20% orange juice; another is 5% orange juice. How many liters of each should be mixed together in order to get 15L that is 6% orange juice?
Answer by mananth(16946)

percent ---------------- quantity
Juice type I 20.00% ---------------- x liters
juice type II 5.00% ------ 15 - x liters
Mixture 6.00% ---------------- 15
Total 15 liters
20.00% x + 5.00% ( 15 - x ) = 6.00% * 15
20 x + 5 ( 15 - x ) = 90
20 x + 75 - 5 x = 90
20 x - 5 x = 90 - -75
15 x = 15
/ 15
x = 1 liters 20.00% Juice type I
14 liters 5.00% juice type II

m.ananth@hotmail.ca